OpenStudy (anonymous):

When the shuttle bus comes to a sudden stop to avoid hitting a dog, it decelerates uniformly at 3.8 m/s2 as it slows from 9.8 m/s to 0 m/s. Find the time interval of acceleration for the bus. Answer in units of s

OpenStudy (anonymous):

Think about your units. You have the deceleration wich is in m/s^2 and the velocity in m/s. If you think about it like that you simly need to divide \[m/s \div m/s^2\] to get your answer in seconds. Hence 9.8/3.8 = 2.58 seconds

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

A car traveling at 6.7m/s accelerates 3.1m/s2 to reach a speed of 11.0 m/s. How long does it take for this acceleration to occur? Answer in units of s

OpenStudy (anonymous):

You need to increase your speed frpm 6.7 m/s to 11m/s. Try to think that your 6.7m/s is your constant speed and you could equalise it to zero since it doen't matter. This means that you have to accelerate 11-6.7 = 4.3m/s Again you need to pay attention the the units. Accelaration has m/s^2 and speed m/s. again if you divide (m/s)/(m/s^2) =(ms^2)/(ms) = second Hence 4.3/3.1 = 1.387 seconds

OpenStudy (anonymous):

Turner’s treadmill starts with a velocity of −2.7 m/s and speeds up at regular intervals during a half-hour workout.After 33 min, the treadmill has a velocity of −7.1 m/s. What is the average acceleration of the treadmill during this period? Answer in units of m/s2

OpenStudy (anonymous):

This time you have to calculate the accelaration. Again the acceleration units is m/s^2. this means you need to divide the velocity you have with the time given. However, you need to convert the minutes in to seconds in order to be concistent. Thus 7.1-2.7 = 4.4m/s divide by 33minutes * 60 to convert to seconds = 1980 secoonds\Hence your answer is 4.4/1980 = 0.0022m/s2